# Sample Complexity of Distributionally Robust Average-Reward Reinforcement Learning

### v1

@@ G.1 DR-DMDP under $f_{k}$ Uncertainty Set
$\leq 12\sup_{Q\in\mathcal{P}}t_{\mathrm{minorize}}(Q_{\pi})\sqrt{\frac{8}{n\mathfrak{p}_{\wedge}}\log\frac{2|\mathbf{S}|^{2}|\mathbf{A}|}{\beta}}$
$\leq 24t_{\mathrm{minorize}}\sqrt{\frac{8}{n\mathfrak{p}_{\wedge}}\log\frac{2|\mathbf{S}|^{2}|\mathbf{A}|}{\beta}}.$
@@ G.2 Reduction to DR-DMDP Approach under $f_{k}$ Uncertainty Set
$0\leq g_{\mathcal{P}}^{*}-g_{\mathcal{P}}^{\widehat{\pi}^{*}}\leq$ $120t_{\mathrm{minorize}}\sqrt{\frac{2}{n\mathfrak{p}_{\wedge}}\log\frac{8|\mathbf{S}|^{2}|\mathbf{A}|}{\beta}}$ (G.4)
$\left\|\frac{V_{\widehat{\mathcal{P}}}^{*}}{\sqrt{n}}-g_{\mathcal{P}}^{*}\right\|_{\infty}\leq$ $60t_{\mathrm{minorize}}\sqrt{\frac{2}{n\mathfrak{p}_{\wedge}}\log\frac{2|\mathbf{S}|^{2}|\mathbf{A}|}{\beta}}$
@@
$\max_{\pi\in\Pi}\frac{1}{\sqrt{n}}\left\|V_{\widehat{\mathcal{P}}}^{\pi}-V_{\mathcal{P}}^{\pi}\right\|_{\infty}\leq$ $\sup_{Q\in\mathcal{P}}\max_{\pi\in\Pi}24t_{\mathrm{minorize}}(Q_{\pi})\sqrt{\frac{8}{n\mathfrak{p}_{\wedge}}\log\frac{2|\mathbf{S}|^{2}|\mathbf{A}|}{\beta}}$ (G.6)
$\leq$ $48t_{\mathrm{minorize}}\sqrt{\frac{8}{n\mathfrak{p}_{\wedge}}\log\frac{2|\mathbf{S}|^{2}|\mathbf{A}|}{\beta}}$
@@
$0\leq g_{\mathcal{P}}^{*}-g_{\mathcal{P}}^{\widehat{\pi}^{*}}\leq$ $\frac{36t_{\mathrm{minorize}}}{\sqrt{n}}+48t_{\mathrm{minorize}}\sqrt{\frac{8}{n\mathfrak{p}_{\wedge}}\log\frac{2|\mathbf{S}|^{2}|\mathbf{A}|}{\beta}}$ (G.7)
@@
$\stackrel{{\scriptstyle(i)}}{{\leq}}$ $120t_{\mathrm{minorize}}\sqrt{\frac{2}{n\mathfrak{p}_{\wedge}}\log\frac{2|\mathbf{S}|^{2}|\mathbf{A}|}{\beta}}$
$\left\|\frac{V_{\widehat{\mathcal{P}}}^{*}}{\sqrt{n}}-g_{\mathcal{P}}^{*}\right\|_{\infty}\stackrel{{\scriptstyle(ii)}}{{\leq}}$ $60t_{\mathrm{minorize}}\sqrt{\frac{2}{n\mathfrak{p}_{\wedge}}\log\frac{2|\mathbf{S}|^{2}|\mathbf{A}|}{\beta}}$
@@ G.3 Anchored DR-AMDP Approach under $f_{k}$ Uncertainty Set
$0\leq g_{\mathcal{P}}^{*}-g_{\mathcal{P}}^{\widehat{\pi}^{*}}\leq$ $120t_{\mathrm{minorize}}\sqrt{\frac{2}{n\mathfrak{p}_{\wedge}}\log\frac{2|\mathbf{S}|^{2}|\mathbf{A}|}{\beta}}$ (G.8)
$\left\|g_{\underline{\widehat{\mathcal{P}}}}^{*}-g_{\mathcal{P}}^{*}\right\|_{\infty}\leq$ $60t_{\mathrm{minorize}}\sqrt{\frac{2}{n\mathfrak{p}_{\wedge}}\log\frac{2|\mathbf{S}|^{2}|\mathbf{A}|}{\beta}}.$
@@
$\left\|g_{\underline{\widehat{\mathcal{P}}}}^{\pi}-g_{\mathcal{P}}^{\pi}\right\|_{\infty}\leq$ $18\frac{t_{\mathrm{minorize}}}{\sqrt{n}}+24t_{\mathrm{minorize}}\sqrt{\frac{8}{n\mathfrak{p}_{\wedge}}\log\frac{2|\mathbf{S}|^{2}|\mathbf{A}|}{\beta}}$ (G.12)
@@
$\leq$ $60t_{\mathrm{minorize}}\sqrt{\frac{2}{n\mathfrak{p}_{\wedge}}\log\frac{2|\mathbf{S}|^{2}|\mathbf{A}|}{\beta}}.$
@@
$0\leq g_{\mathcal{P}}^{*}-g_{\mathcal{P}}^{\widehat{\pi}^{*}}\leq$ $120t_{\mathrm{minorize}}\sqrt{\frac{2}{n\mathfrak{p}_{\wedge}}\log\frac{2|\mathbf{S}|^{2}|\mathbf{A}|}{\beta}}$ (G.13)
$\left\|g_{\underline{\widehat{\mathcal{P}}}}^{*}-g_{\mathcal{P}}^{*}\right\|_{\infty}\leq$ $60t_{\mathrm{minorize}}\sqrt{\frac{2}{n\mathfrak{p}_{\wedge}}\log\frac{2|\mathbf{S}|^{2}|\mathbf{A}|}{\beta}}.$

### v2

@@ G.1 DR-DMDP under $f_{k}$ Uncertainty Set
[…]
@@ G.2 Reduction to DR-DMDP Approach under $f_{k}$ Uncertainty Set
[…]
@@
[…]
@@
[…]
@@
[…]
@@ G.3 Anchored DR-AMDP Approach under $f_{k}$ Uncertainty Set
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Appendix H Additional Experiment
To further expand on our results in Section 4, we provide additional experiments on baseline comparison and large-scale MDPs. First, we include comparisons with DR RVI Q-learning (Wang et al., 2023e, ) on Hard MDP Instance 5.1 as baseline. Table 2 and Figure 3 shows the error performance for DR RVI Q-learning and the two algorithms for $\mathfrak{p}_{\wedge}=0.9$ and $\delta=0.1$. They demonstrated that the error levels of our two algorithms are comparable and significantly outperform the previous baseline.
Table 2: Performance comparison with DR RVI Q-learning.
Sample 10 32 100 316 1000 3162 10000 31622 100000
DR RVI Q-learning (Wang et al., 2023e, ) 1.84e-1 7.36e-2 7.47e-2 6.20e-2 5.52e-2 4.46e-2 3.60e-2 3.08e-2 2.61e-2
Reduction to DMDP 1.21e-1 5.95e-2 5.35e-2 2.74e-2 1.39e-2 6.65e-3 3.49e-3 3.21e-3 1.17e-3
Anchored DR-AMDP 1.67e-1 6.52e-2 6.26e-2 2.90e-2 1.16e-2 7.51e-3 3.27e-3 2.37e-3 1.23e-3
Figure 3: KL-divergence case for Algorithm 2 and 3
Further, to demonstrate the capability of our framework in solving large-scale MDPs, we consider a large-scale MDP with $20$ states and $30$ actions, as in Wang et al., 2023e . The nominal transition distribution is specified by $p_{s,a}\sim\mathcal{N}(1,\sigma_{s,a})$, where $\sigma_{s,a}\sim\textbf{Uniform}[0,100]$, followed by normalization. We then choose the uncertainty size $\delta=0.4$ to introduce stronger perturbations, following the setting in Wang et al., 2023e under the KL-divergence. Note that although $\delta=0.4$ violates Assumption 2, the slope of the linear regression of our results on the logarithmic scale remains very close to 1/2-1/2, further supporting our theoretical guarantees. This observation empirically validates the theoretical results established in our theorems.
Further, to demonstrate the capability of our framework in solving large-scale MDP instances, we consider a large-scale MDP with $20$ states and $30$ actions, as in Wang et al., 2023e . Specifically, Algorithm 2 and Algorithm 3 are evaluated on two distinct large-scale instances, respectively, to verify their effectiveness. The nominal transition distribution is specified by $p_{s,a}\sim\mathcal{N}(1,\sigma_{s,a})$, where $\sigma_{s,a}\sim\textbf{Uniform}[0,100]$, followed by normalization. We then choose the uncertainty size $\delta=0.4$ to introduce stronger perturbations, following the setting in Wang et al., 2023e under the KL-divergence. Note that although $\delta=0.4$ violates Assumption 2, the slope of the linear regression of our results on the logarithmic scale remains very close to 1/2-1/2, further supporting our theoretical guarantees. This observation empirically validates the theoretical results established in our theorems.
Table 3: Performance on large-scale MDP across different sample sizes.
Sample 10 32 100 316 1000 3162 10000 31622 100000
Reduction to DMDP 1.04e-1 6.54e-2 3.07e-2 2.46e-2 8.28e-3 4.82e-3 4.66e-3 3.12e-3 1.12e-3
Anchored DR-AMDP 6.55e-2 3.72e-2 2.69e-2 4.79e-3 3.73e-3 1.15e-3 1.33e-3 9.42e-4 5.61e-4
a Algorithm 2 for large-scale MDP.
b Algorithm 3 for large-scale MDP.